OpenStudy (anonymous):

What is the solution to the equation 9^(x + 1) = 27 ?

OpenStudy (amistre64):

can we rewrite 27 as a base 9?

OpenStudy (amistre64):

we can prolly do a base 3 on them both tho

OpenStudy (amistre64):

9^3 ??

OpenStudy (anonymous):

snaffoooo

OpenStudy (amistre64):

3^3 = 27 :)

OpenStudy (amistre64):

log(3) the sides

OpenStudy (anonymous):

:) I erased it to avoid further confusion

OpenStudy (amistre64):

log3(9^(x + 1)) = log3(27) (x+1) log3(9) = 3 x+1 = 3/2 x = 3/2 - 1 maybe?

OpenStudy (anonymous):

x = –0.5 x = 2.5 x = 0.5 x = –3.5

OpenStudy (amistre64):

i dont have a decimater

OpenStudy (anonymous):

\[3^{2(x+1)}=3^3\] 2(x+1)=3 2x+2=3 2x=1 x=0.5

OpenStudy (anonymous):

Well out of those 4, is it 0.5?

OpenStudy (amistre64):

id say it is :)

OpenStudy (anonymous):

Thanks! c: